Status Report

NASA Space Station On-Orbit Status 28 Jun 2004

By SpaceRef Editor
June 29, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 28 Jun 2004

All ISS systems continue to function nominally except those noted previously or below.   Underway: Week 9 of Increment 9.  Day 1304 of permanent human station residency, and Day 2048 since first ISS launch (FGB).

The crew continued preparations for the Orlan EVA (Take 2) on Wednesday, 6/30, readying the suits and their consumable/replaceable ORU elements.  Hatch open for the 6-hr. spacewalk is scheduled for 5:40pm EDT.   [In a repeat of last week’s pre-EVA ops, Padalka and Fincke performed leak checks and valve functionality tests on the suits and their BSS interface units, replaced oxygen bottles and set up communications, with checks of Orlan and BSS telemetry, voice, and biomedical parameter transmission.  Gennady also temporarily disconnected the SM’s BITS2-12 telemetry system to allow deactivation and teardown of the ZU-S battery charger.  Orlan ORUs are LiOH canisters (LP-9), primary & backup oxygen tanks (BK-3), moisture collectors, feedwater filters (FOR), CO2 measuring unit (IK) filter, filtration & separation units (BOS), and the newly charged 825M1 storage batteries.]

The crew had time reserved for another review of the EVA timeline and a checkup on the already packed EVA tool & gear bundle, including the spare RPCM.

Also in preparation for the spacewalk, FE/SO Fincke terminated the re-charging of the U.S. NiMH (nickel metal hydride) batteries for the helmet lights to be used with the Orlans during darkness periods.

In the Lab module, Mike powered up the HRF GASMAP (Human Research Facility/Gas Analyzer System for Metabolic Analysis Physiology) equipment for its periodic functional checkout.   [Today’s test run was a once-every-increment modification of the regular monthly health check that Mike performed previously.]

Gennady conducted the periodic (monthly) functional closure test of the Vozdukh CO2 removal system’s emergency vacuum valves (AVK, last time done: 5/24).   [The AVKs are critical because they close the Vozdukh’s vacuum access lines in the event of a malfunction in the regular vacuum valves (BVK) or a depressurization in the Vozdukh valve panel (BOA).  Access to vacuum is required to vent carbon dioxide (CO2) during the regeneration of the absorbent cartridges (PP).  During nominal operation, the AVK valves remain open.]

The crew completed the periodic station air sampling.  The CDR first tested for CO (carbon monoxide) and NH3 (ammonia) levels in SM and DC-1 by collecting samples with the IPD Draeger tubes; then took readings for air constituents in SM and FGB with the AK-1M sampler.  The FE meanwhile “grabbed” air samples with the GSC (grab sample container) at the center of the SM, then used the new Dual Sorbent Tube (DST), instead of the old SSAS (Solid Sorbent Air Sampler), to gather samples in the center of the Lab and SM.
The Elektron O2 generator continues to operate nominally, after the brief shutdown of both micropumps over the weekend due to air bubbles in the BZh liquid unit causing impeller cavitation.   [Without the Elektron, O2 quantities on board from the U.S. Airlock HPTs (high-pressure tanks), SFOGs (solid fuel oxygen generators) and Progress tanks would currently sustain the crew for an estimated 97 days.]

TsUP/Moscow is planning to run a brief test with the Progress 14P MCS (motion control system) tonight, to verify a new propellant-saving procedure for USOS/RS attitude control handovers and maneuvers.  The technique, which works only in LVLH (local vertical/local horizontal) attitude, will be tested from 7:45pm-11:20pm EDT, i.e., during crew sleep.   [The new prop saving technique eliminates two thruster maneuvers (to MPA/minimum propellant attitude and back to TEA/torque equilibrium attitude), with the CMG gimbals being positioned more efficiently to compensate for the MPA elimination.  In computer studies the new procedure has used ~4 kg props for the 3-hr period under RS control, instead of ~18 kg required without it.]

Both crewmembers completed their daily 2.5-hr. physical exercise program on TVIS, RED exerciser, and VELO cycle with load trainer.

Today’s CEO (Crew Earth Observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Internal waves, Aegean Sea (looking left towards the glint point which may illuminate much of the width of the sea.  The features are small enough to suggest a longer focal length), Athens, Greece (looking a touch right for the city, which may have required more than one image), Algiers, Algeria (nadir pass, on the coast), Impact Crater Field, Libya (newly discovered impact field.  Although the impacts occurred hundreds of millions years ago, the craters have been recently re-exposed by erosion of the cover strata.  Suggested were less detailed images to capture the general area.  Detailed images will be requested later), El Paso, Texas (nadir pass), and Monterrey, Mexico (nadir pass).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. & Russian Segment Status  (as of today, 1:27pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is now functioning again).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 25.6; ppO2 (mmHg) — 157.0; ppCO2 (mmHg) — 4.6.
  • SM Transfer Compartment:  Pressure (mmHg) — 741; temperature (deg C) — 20.2.
  • FGB Cabin:  Pressure (mmHg) — 750; temperature (deg C) — 20.1.
  • Node:  Pressure (mmHg) — 745.80; temperature (deg C) — 23.6 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 747.85; temperature (deg C) — 24.6; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 747.05; temperature (deg C) — 22.8; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) — n/a
  • PMA-2:  Shell heater temp (deg C) — n/a.

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Dual angle/”blind” mode, non solar-tracking, biased for drag reduction).
  • SM batteries:  All batteries (8) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #5 is off line; all other batteries (5) are on line in “Partial Charge” mode.

Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3963 kg (8737 lb) as of 6/24/04;  [SM(552) + FGB(2772) + Progress M(639)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 2 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management, until 6/28, following the EVA.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:49am EDT [= epoch]):

  • Mean altitude — 360.6 km
  • Apogee — 364.2 km
  • Perigee — 356.9 km
  • Period — 91.7 min.
  • Inclination (to Equator) —  51.6319 deg
  • Eccentricity — 0.0005434
  • Solar Beta Angle — 20.3 deg
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss in last 24 hours — 85 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32009

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.